OpenStudy (anonymous):

HELP!!MEDALS!!FAN!! Mandy's Candies is most famous for their homemade fudge and homemade toffee. They make an average of 146 pounds of fudge and toffee combined each month. They sell a pound of fudge for $9.56 and a pound of toffee for $7.21. Mandy's averages $1,212.46 each month in fudge and toffee sales. On average, how many pounds each of fudge and toffee does Mandy's Candies sell in one month? 71 fudge; 75 toffee 68 fudge; 78 toffee 74 fudge; 72 toffee 57 fudge; 89 toffee

OpenStudy (anonymous):

@ketty @Yuba @RockStar02 @csscgirl

OpenStudy (anonymous):

if you solve the system of equations, you can find the answer... here is the system: 9.56x+7.21y=1212.46 x+y=1752... someone needs to check me, because i may not have it right, but i am pretty sure if you solve that system, you'll get the answer.

OpenStudy (anonymous):

so the answer....

OpenStudy (anonymous):

b

OpenStudy (anonymous):

the second equation from the system of equations i gave you was wrong... it should have been x+y=146.

OpenStudy (anonymous):

dont worry my session timed out before i could put any answers